OpenStudy (ujjwal):

Find a necessary and sufficient condition on \(a\in \mathbb{R}\) for a quadratic equation \(x^2+2(a+3)x-a+3=0\) to have roots such that 1)both of them are positive 2)both of them are negative

OpenStudy (anonymous):

Use "\in", not "\epsilon". The roots of the quadratic are the solutions to \[x^2+2(a+3)\color{red}x-a+3\] Am I right to assume there's supposed to be an \(x\) there?

OpenStudy (ujjwal):

lol. yeah.. I missed x

OpenStudy (anonymous):

I left out an \(=0\) at the end of that equation... Anyway, the solutions to this equation, given by the quadratic formula, are \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{-2(a+3)\pm\sqrt{(2(a+3))^2-4(-a+3)}}{2}\] Which values of \(a\) give you both positive/both negative roots?

OpenStudy (anonymous):

$$x^2+2(a+3)x-(a-3)$$notice with roots \(u,v\) we have:$$-2(a+3)=u+v\\-(a-3)=uv$$if both \(u,v>0\) it follows that \(u+v>0\) and \(uv>0\), which suggests \(a+3<0\Leftrightarrow a<-3\) and \(a-3<0\Leftrightarrow a<3\). note \((a<-3)\land (a<3)=a<-3\) if instead \(u,v<0\) we have \(u+v<0\) yet \(uv>0\) suggesting \(a+3>0\Leftrightarrow a>-3\) and also \(a-3<0\Leftrightarrow a<3\) hence \(-3<a<3\)

OpenStudy (anonymous):

this answer is incomplete, though; you also need that two real roots exist, which requires showing the discriminant is positive i.e.$$(2(a+3))^2-4(-(a-3))>0\\4(a^2+6a+9)+4(a-3)>0\\4(a^2+7a+6)>0\\a^2+7a+6>0\\(a+1)(a+6)>0$$

OpenStudy (anonymous):

clearly for \(a>-1\) or \(a<-6\) the above is satisfied hence our inequalities are actually:$$(a<-3)\land(a<-6)\implies a<-6\\(-3<a<3)\land(a>-1)\implies -1<a<3$$

OpenStudy (ujjwal):

Thank you!! :)

OpenStudy (anonymous):

@SithsAndGiggles finish your approach... I want to see if it's easier :-p

OpenStudy (anonymous):

@oldrin.bataku, whether it is or not, I'm not sure, but it all comes down to solving a bunch of inequalities like with your method. Like you said, the discriminant must be positive, so you have \[4(a+3)^2+4a-12>0~~\Rightarrow~~a>-1~\text{ and }~a<-6\] For both roots to be positive, you must have \[-2(a+3)+\sqrt{(2(a+3))^2-4(-a+3)}>0~~~~~~~~~~~~~~~~a>3~~\text{and}~~a\le-6\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Rightarrow\\ -2(a+3)-\sqrt{(2(a+3))^2-4(-a+3)}>0~~~~~~~~~~~~~~~a\le-6 \] (< 0 for both to be negative roots, which gives \(-1\le a<3\) for the added root and \(a\ge-1\) for the subtracted root.) There seems to be a disparity with our inequality signs, but I used WA to find the solutions quickly.